A certain group of test subjects had pulse rates with a mean of 81.9 beats per minute and a standard deviation of 11.8 beats per minute. Would it be unusual for one of the test subjects to have a pulse rate of 125.5 beats per minute?

Minimum "usual" value?

Maximum "usual" value?

Is 125.5 beats per minute an unusual pulse rate?
Answer #1

Mean = 81.9 beats per minute

Standard deviation = 11.8 beats per minute

A value are considered usual if it lies within 2 standard deviation of mean

Minimum usual value = Mean - 2 x standard deviation

= 81.9 - 2x11.8

= 58.3 beats per minute

Maximum usual value = Mean + 2 x standard deviation

= 81.9 + 2x11.8

= 105.5 beats per minute

125.5 is larger than the maximum usual value.

So, 125.5 beats per minute is an unusual pulse rate